# Scalable and consistent embedding of probability measures into Hilbert spaces via measure quantization

### v1

@@ Appendix A Proofs of the main results
$W_{2}^{2}(\bar{\nu}^{(i)}_{K},\bar{\nu}^{(j)}_{K})=\max_{(\alpha,\beta)\in\Phi}\ \sum_{k=1}^{K}a_{k}^{(i)}\alpha_{k}+\sum_{k=1}^{K}a_{k}^{(j)}\beta_{k}$ (32)
where $\Phi:=\{(\alpha,\beta)\in\mathbb{R}^{K}\times\mathbb{R}^{K}\ \mbox{such that for all}\ 1\leq k,l\leq K,\alpha_{k}+\beta_{l}\leq C_{kl}\}$. Let $\alpha^{ij},\beta^{ij}\in\mathbb{R}^{K}$ be optimal Kantorovich potentials for $\bar{\nu}^{(i)}_{K}$ and $\bar{\nu}^{(j)}_{K}$ in (32). We define the piecewise constant function $f^{ij}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ such that $x\mapsto\alpha^{ij}_{k}$ when $x\in\overset{\circ}{V_{k}}$, where $\overset{\circ}{V_{k}}$ denotes the open interior of the Voronoï cell $V_{k}$. Similarly, $g^{ij}:y\mapsto\beta^{ij}_{k}$ when $y\in V_{k}$. Then, thanks to the absolute continuity of the $\mu^{(i)}$’s, one can write:
@@
$=\sum\limits_{k=1}^{K}\int_{\overset{\circ}{V_{k}}}\alpha^{ij}_{k}\mathrm{d}\mu^{(i)}(x)+\sum\limits_{k=1}^{K}\int_{\overset{\circ}{V_{k}}}\beta^{ij}_{k}\mathrm{d}\mu^{(j)}(y)$
$=\int_{\mathbb{R}^{d}}f^{ij}(x)\mathrm{d}\mu^{(i)}(x)+\int_{\mathbb{R}^{d}}g^{ij}(y)\mathrm{d}\mu^{(j)}(y)$ (33)
We then aim at identifying (33) with a dual formulation for OT between $\mu^{(i)}$ and $\mu^{(j)}$ with respect to some cost function $c:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ to be defined later on, where
@@
$=\sup\limits_{f,g:f(x)+g(y)\leq c(x,y)}\int_{\mathbb{R}^{d}}f(x)\mathrm{d}\mu^{(i)}(x)+\int_{\mathbb{R}^{d}}g(y)\mathrm{d}\mu^{(j)}(y).$ (34)
@@
$\leq 6\max\limits_{k}\mathrm{diam}(V_{k})+3\|x-y\|^{2},$ (35)
where the first inequality is due to the fact that $\alpha^{ij}_{k}$ and $\beta^{ij}_{k^{\prime}}$ are Kantorovich potentials of $W_{2}(\bar{\nu}^{(i)}_{K},\bar{\nu}^{(j)}_{K})$ in (32). The second inequality comes from (30) where $\lambda=1$. Defining the new cost function $c(x,y)=6\max\limits_{k}\mathrm{diam}(V_{k})+3\|x-y\|^{2}$, we thus define
@@
Now, by inequality (35), it follows that $f^{ij}$ and $g^{ij}$ are feasible Kantorovich potentials of $\text{OT}_{c}$ between $\mu^{(i)}$ and $\mu^{(j)}$ in (34). Hence, we finally obtain from (33) that
@@
For the result regarding the within-class variance of the clusters, we have, using the triangle inequality and (30) with $\lambda=1$,
$W_{2}^{2}(\nu^{(i)}_{K},\nu^{(j)}_{K})$ $\leq 3\bigl{(}W_{2}^{2}(\nu^{(i)}_{K},\mu^{(i)})+W_{2}^{2}(\mu^{(i)},\mu^{(j)})+W_{2}^{2}(\mu^{(j)},\nu^{(j)}_{K})\bigr{)}$ (36)
@@
where the last equality follows from Proposition 3.4, which concludes the first item (19) of the proposition. For the second statement on the between-class variance, we rewrite the triangle inequality:
@@
where the last equality follows from Proposition 3.4, which concludes the proof. ∎
@@
Let $\phi:\mathcal{P}(\mathcal{X})\rightarrow\mathcal{H}$ be an embedding of probability measures (supported on an arbitrary set included in $\mathcal{X}\subset\mathbb{R}^{d}$) into a Hilbert space $\mathcal{H}$ equipped with the inner-product $\langle\cdot,\cdot\rangle_{\mathcal{H}}$ and the induced norm $\|\cdot\|_{\mathcal{H}}$. Suppose that $\|\phi\|_{\infty}=\sup\limits_{\mu\in\mathcal{P}(\mathcal{X})}\|\phi(\mu)\|_{\mathcal{H}}<\infty$. Then, we have that:
$\|G^{\phi}_{\mu}-G^{\phi}_{\nu_{K}}\|_{F}^{2}\leq C\sum\limits_{i=1}^{N}\|\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\|^{2}_{\mathcal{H}}$ (37)
where $C$ is a constant depending on $N$ and $\|\phi\|_{\infty}$.
@@
$\|G^{\phi}_{\mu}-G^{\phi}_{\nu_{K}}\|_{F}^{2}=\sum\limits_{i,j=1}^{N}|(G^{\phi}_{\mu}-G^{\phi}_{\nu_{K}})_{ij}|^{2}=\sum\limits_{i,j=1}^{N}|\langle\phi(\mu^{(i)}),\phi(\mu^{(j)})\rangle_{\mathcal{H}}-\langle\phi(\nu^{(i)}_{K}),\phi(\nu^{(j)}_{K})\rangle_{\mathcal{H}}|^{2}$ (38)
@@
$\langle\phi(\mu^{(i)}),\phi(\mu^{(j)})\rangle_{\mathcal{H}}$ $-\langle\phi(\nu^{(i)}_{K}),\phi(\nu^{(j)}_{K})\rangle_{\mathcal{H}}=\langle\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K}),\phi(\mu^{(j)})\rangle_{\mathcal{H}}+\langle\phi(\nu^{(i)}_{K}),\phi(\mu^{(j)})\rangle_{\mathcal{H}}$
$\qquad-\langle\phi(\nu^{(i)}_{K}),\phi(\nu^{(j)}_{K})\rangle_{\mathcal{H}}$
$=\langle\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K}),\phi(\mu^{(j)})-\phi(\nu^{(j)}_{K})\rangle_{\mathcal{H}}+\langle\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K}),\phi(\nu^{(j)}_{K})\rangle_{\mathcal{H}}$
$\quad+\langle\phi(\nu^{(i)}_{K}),\phi(\mu^{(j)})\rangle_{\mathcal{H}}-\langle\phi(\nu^{(i)}_{K}),\phi(\nu^{(j)}_{K})\rangle_{\mathcal{H}}$
$=\langle\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K}),\phi(\mu^{(j)})-\phi(\nu^{(j)}_{K})\rangle_{\mathcal{H}}+\langle\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K}),\phi(\nu^{(j)}_{K})\rangle_{\mathcal{H}}$
$\quad+\langle\phi(\mu^{(j}))-\phi(\nu^{(j)}_{K}),\phi(\nu^{(i)}_{K})\rangle_{\mathcal{H}}.$
Injecting this equality in (38) yields:
@@
$\leq 3\sum\limits_{i,j=1}^{N}\Bigl{\|}\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\Bigr{\|}_{\mathcal{H}}^{2}\Bigl{\|}\phi(\mu^{(j)})-\phi(\nu^{(j)}_{K})\Bigr{\|}_{\mathcal{H}}^{2}$ (39)
@@
$\qquad+\sum\limits_{i,j=1}^{N}\Bigl{\|}\phi(\mu^{(j}))-\phi(\nu^{(j)}_{K})\Bigr{\|}_{\mathcal{H}}^{2}\Bigl{\|}\phi(\nu^{(i)}_{K})\Bigr{\|}_{\mathcal{H}}^{2}$
@@
$\leq\sum\limits_{i=1}^{N}\bigl{\|}\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\bigr{\|}^{2}_{\mathcal{H}}\Bigl{(}3\sum\limits_{i=1}^{N}\bigl{\|}\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\bigr{\|}^{2}_{\mathcal{H}}+6\sum\limits_{j=1}^{N}\bigl{\|}\phi(\nu^{(j)}_{K})\bigr{\|}^{2}_{\mathcal{H}}\Bigr{)}$ (40)
where (39) is due to Cauchy-Schwarz. By hypothesis, we have $\|\phi\|_{\infty}=\sup\limits_{\mu\in\mathcal{P}(\mathcal{X})}\|\phi(\mu)\|_{\mathcal{H}}=M_{\phi}<\infty$. Therefore, (40) yields:
$\|G^{\phi}_{\mu}-G^{\phi}_{\nu_{K}}\|_{F}^{2}$ $\leq\sum\limits_{i=1}^{N}\bigl{\|}\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\bigr{\|}^{2}_{\mathcal{H}}\Bigl{(}3\sum\limits_{i=1}^{N}M_{\phi}^{2}+6\sum\limits_{j=1}^{N}M_{\phi}^{2}\Bigr{)}$
$=9NM_{\phi}^{2}\sum\limits_{i=1}^{N}\bigl{\|}\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\bigr{\|}^{2}_{\mathcal{H}}.$
Choosing $C:=9NM_{\phi}^{2}$ concludes the proof.
@@
(ii) We first prove inequality (22) for the KME embedding $\phi:\sigma\in\mathcal{P}(\mathcal{X})\mapsto\int k(x,\cdot)\mathrm{d}\sigma(x)\in\mathcal{H}$. By hypothesis, the kernel $k$ is bounded by a constant $M_{k}$. Therefore for any $\mu\in\mathcal{P}(\mathcal{X})$, we have $\|\phi(\mu)\|_{\mathcal{H}}^{2}=\langle\int k(x,\cdot)\mathrm{d}\mu(x),\int k(y,\cdot)\mathrm{d}\mu(y)\rangle_{\mathcal{H}}=\iint k(x,y)\mathrm{d}\mu(x)\mathrm{d}\mu(y)\leq M_{k}$. By Lemma A.2, we then have
@@
with $C_{N,k}:=9NM_{k}$. Using the property $\mathrm{MMD}\leq W_{1}\leq W_{2}$, we obtain that
$\frac{1}{N}\|G_{\mu}^{\mathrm{KME}}-G_{\nu_{K}}^{\mathrm{KME}}\|_{F}^{2}$ $\leq\frac{C_{N,k}}{N}\sum\limits_{i=1}^{N}\|\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\|^{2}_{\mathcal{H}}=\frac{C_{N,k}}{N}\sum\limits_{i=1}^{N}\mathrm{MMD}^{2}(\mu^{(i)},\nu^{(i)}_{K})$
$\leq\frac{C_{N,k}}{N}\sum\limits_{i=1}^{N}W_{2}^{2}(\mu^{(i)},\nu^{(i)}_{K})$
$=C_{N,k}\ \varepsilon_{K}.$
(i) In order to prove inequality (21) for the LOT embedding $\phi:\mu\in\mathcal{P}(\mathcal{X})\mapsto T_{\rho}^{\mu}-\mathrm{id}\in L^{2}(\rho)$ with a.c. reference measure $\rho$, we first note that for any measure $\mu\in\mathcal{P}(\mathcal{X})$, $\|\phi(\mu)\|^{2}_{L^{2}(\rho)}=\int|T_{\rho}^{\mu}(x)-x|^{2}\mathrm{d}\rho(x)\leq\int\mathrm{diam}(\mathcal{X})^{2}\mathrm{d}\rho(x)=\mathrm{diam}(\mathcal{X})^{2}$. Then we can apply Lemma A.2 with constant $C:=C_{N,\mathcal{X}}=9N\mathrm{diam}(\mathcal{X})^{2}$ in (37) and finally obtain
$\frac{1}{N}\|G_{\mu}^{\mathrm{LOT}}-G_{\nu_{K}}^{\mathrm{LOT}}\|_{F}^{2}$ $\leq\frac{C_{N,\mathcal{X}}}{N}\sum\limits_{i=1}^{N}\|T_{\rho}^{\mu^{(i)}}-T_{\rho}^{\nu^{(i)}_{K}}\|_{L^{2}(\rho)}^{2}.$ (41)
@@
$\|T_{\rho}^{\mu}-T_{\rho}^{\nu}\|_{L^{2}(\rho)}\leq 2\sqrt{\mathrm{diam}(\mathcal{X})L}W_{1}(\mu,\nu)^{1/2}.$ (42)
Finally, assembling the inequalities (41) and (42), we obtain
@@
$=4C_{N,\mathcal{X}}\mathrm{diam}(\mathcal{X})L\frac{1}{N}\sqrt{\Bigl{(}\sum\limits_{i=1}^{N}W_{2}(\mu^{(i)},\nu^{(i)}_{K})\Bigr{)}^{2}}$
@@
$=4C_{N,\mathcal{X}}\mathrm{diam}(\mathcal{X})L\frac{1}{N}\sqrt{N\cdot N\varepsilon_{K}(\overline{\mu})}$
$=4C_{N,\mathcal{X}}\mathrm{diam}(\mathcal{X})L\cdot\sqrt{\varepsilon_{K}(\overline{\mu})}$
$=C_{N,\mathcal{X},L}\sqrt{\varepsilon_{K}(\overline{\mu})},$

### v3

@@ Appendix A Proofs of the main results
$W_{2}^{2}(\bar{\nu}^{(i)}_{K},\bar{\nu}^{(j)}_{K})=\max_{(\alpha,\beta)\in\Phi}\ \sum_{k=1}^{K}a_{k}^{(i)}\alpha_{k}+\sum_{k=1}^{K}a_{k}^{(j)}\beta_{k}$ (A.8)
where $\Phi:=\{(\alpha,\beta)\in\mathbb{R}^{K}\times\mathbb{R}^{K}\ \mbox{such that for all}\ 1\leq k,l\leq K,\alpha_{k}+\beta_{l}\leq\|\bar{x}_{k}-\bar{x}_{l}\|^{2}\}$. Let $\alpha^{ij},\beta^{ij}\in\mathbb{R}^{K}$ be optimal Kantorovich potentials for $\bar{\nu}^{(i)}_{K}$ and $\bar{\nu}^{(j)}_{K}$ in (A.8). We define the piecewise constant function $f^{ij}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ such that $x\mapsto\alpha^{ij}_{k}$ when $x\in\textrm{int}(V_{k})$, where $\textrm{int}(V_{k})$ denotes the open interior of the Voronoï cell $V_{k}$. Similarly, $g^{ij}:y\mapsto\beta^{ij}_{k}$ when $y\in V_{k}$. Then, thanks to the absolute continuity of the $\mu^{(i)}$’s, one can write:
@@
$=\sum\limits_{k=1}^{K}\int_{\textrm{int}(V_{k})}\alpha^{ij}_{k}\mathrm{d}\mu^{(i)}(x)+\sum\limits_{k=1}^{K}\int_{\textrm{int}(V_{k})}\beta^{ij}_{k}\mathrm{d}\mu^{(j)}(y)$
$=\int_{\mathbb{R}^{d}}f^{ij}(x)\mathrm{d}\mu^{(i)}(x)+\int_{\mathbb{R}^{d}}g^{ij}(y)\mathrm{d}\mu^{(j)}(y)$ (A.9)
We then aim at identifying (A.9) with a dual formulation for OT between $\mu^{(i)}$ and $\mu^{(j)}$ with respect to some cost function $c:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ to be defined later on, where
@@
$=\sup\limits_{f,g:f(x)+g(y)\leq c(x,y)}\int_{\mathbb{R}^{d}}f(x)\mathrm{d}\mu^{(i)}(x)+\int_{\mathbb{R}^{d}}g(y)\mathrm{d}\mu^{(j)}(y).$ (A.10)
@@
$\leq 6\max\limits_{k}\mathrm{diam}(V_{k})+3\|x-y\|^{2},$ (A.11)
where the first inequality is due to the fact that $\alpha^{ij}_{k}$ and $\beta^{ij}_{k^{\prime}}$ are Kantorovich potentials of $W_{2}(\bar{\nu}^{(i)}_{K},\bar{\nu}^{(j)}_{K})$ in (A.8). The second inequality comes from (A.6) where $\lambda=1$. Defining the new cost function $c(x,y)=6\max\limits_{k}\mathrm{diam}(V_{k})+3\|x-y\|^{2}$, we thus define
@@
Now, by inequality (A.11), it follows that $f^{ij}$ and $g^{ij}$ are feasible Kantorovich potentials of $\text{OT}_{c}$ between $\mu^{(i)}$ and $\mu^{(j)}$ in (A.10). Hence, we finally obtain from (A.9) that
@@
For the result regarding the within-class variance of the clusters, we have, using the triangle inequality and (A.6) with $\lambda=1$,
$W_{2}^{2}(\nu^{(i)}_{K},\nu^{(j)}_{K})\leq 3\bigl{(}W_{2}^{2}(\nu^{(i)}_{K},\mu^{(i)})+W_{2}^{2}(\mu^{(i)},\mu^{(j)})+W_{2}^{2}(\mu^{(j)},\nu^{(j)}_{K})\bigr{)}$
@@
where the last equality follows from Theorem 3.4, which concludes the first item (4.1) of the proposition. For the second statement on the between-class variance, we rewrite the triangle inequality:
@@
where the last equality follows from Theorem 3.4, which concludes the proof. ∎
@@
Let $\phi:\mathcal{P}(\mathcal{X})\rightarrow\mathcal{H}$ be an embedding of probability measures (supported on an arbitrary set included in $\mathcal{X}\subset\mathbb{R}^{d}$) into a Hilbert space $\mathcal{H}$ equipped with the inner-product $\langle\cdot,\cdot\rangle_{\mathcal{H}}$ and the induced norm $\|\cdot\|_{\mathcal{H}}$. Suppose that $\|\phi\|_{\infty}=:M_{\phi}<\infty$. Then, we have that:
$\|G^{\phi}_{\mu}-G^{\phi}_{\nu_{K}}\|_{F}^{2}\leq C\sum\limits_{i=1}^{N}\|\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\|^{2}_{\mathcal{H}}$ (A.12)
where $C$ is a constant depending on $N$ and $M_{\phi}$.
@@
$\|G^{\phi}_{\mu}-G^{\phi}_{\nu_{K}}\|_{F}^{2}=\sum\limits_{i,j=1}^{N}|(G^{\phi}_{\mu}-G^{\phi}_{\nu_{K}})_{ij}|^{2}=\sum\limits_{i,j=1}^{N}|\langle\phi(\mu^{(i)}),\phi(\mu^{(j)})\rangle_{\mathcal{H}}-\langle\phi(\nu^{(i)}_{K}),\phi(\nu^{(j)}_{K})\rangle_{\mathcal{H}}|^{2}$ (A.13)
@@
$\langle\phi(\mu^{(i)}),\phi(\mu^{(j)})\rangle_{\mathcal{H}}-\langle\phi(\nu^{(i)}_{K}),\phi(\nu^{(j)}_{K})\rangle_{\mathcal{H}}$
$=\langle\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K}),\phi(\mu^{(j)})-\phi(\nu^{(j)}_{K})\rangle_{\mathcal{H}}+\langle\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K}),\phi(\nu^{(j)}_{K})\rangle_{\mathcal{H}}+\langle\phi(\mu^{(j}))-\phi(\nu^{(j)}_{K}),\phi(\nu^{(i)}_{K})\rangle_{\mathcal{H}}.$
Injecting this equality in (A.13) yields:
@@
$\leq 3\sum\limits_{i,j=1}^{N}\Bigl{\|}\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\Bigr{\|}_{\mathcal{H}}^{2}\Bigl{\|}\phi(\mu^{(j)})-\phi(\nu^{(j)}_{K})\Bigr{\|}_{\mathcal{H}}^{2}$ (A.14)
@@
$\qquad+3\sum\limits_{i,j=1}^{N}\Bigl{\|}\phi(\mu^{(j)})-\phi(\nu^{(j)}_{K})\Bigr{\|}_{\mathcal{H}}^{2}\Bigl{\|}\phi(\nu^{(i)}_{K})\Bigr{\|}_{\mathcal{H}}^{2}$
@@
$\leq\sum\limits_{i=1}^{N}\bigl{\|}\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\bigr{\|}^{2}_{\mathcal{H}}\Bigl{(}3\sum\limits_{i=1}^{N}\bigl{\|}\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\bigr{\|}^{2}_{\mathcal{H}}+6\sum\limits_{j=1}^{N}\bigl{\|}\phi(\nu^{(j)}_{K})\bigr{\|}^{2}_{\mathcal{H}}\Bigr{)}$ (A.15)
where (A.14) is due to Cauchy-Schwarz. By hypothesis, we have $\|\phi\|_{\infty}=\sup\limits_{\mu\in\mathcal{P}(\mathcal{X})}\|\phi(\mu)\|_{\mathcal{H}}=M_{\phi}<\infty$. Therefore, (A.15) yields:
$\|G^{\phi}_{\mu}-G^{\phi}_{\nu_{K}}\|_{F}^{2}$ $\leq\sum\limits_{i=1}^{N}\bigl{\|}\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\bigr{\|}^{2}_{\mathcal{H}}\Bigl{(}12\sum\limits_{i=1}^{N}M_{\phi}^{2}+6\sum\limits_{j=1}^{N}M_{\phi}^{2}\Bigr{)}$
$=18NM_{\phi}^{2}\sum\limits_{i=1}^{N}\bigl{\|}\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\bigr{\|}^{2}_{\mathcal{H}}.$
Choosing $C:=18NM_{\phi}^{2}$ concludes the proof.
@@
(ii) We first prove inequality (4.4) for the KME embedding $\phi:\sigma\in\mathcal{P}(\mathcal{X})\mapsto\int k(x,\cdot)\mathrm{d}\sigma(x)\in\mathcal{H}$. By hypothesis, the kernel $k$ is bounded by a constant $M_{k}$. Therefore for any $\mu\in\mathcal{P}(\mathcal{X})$, we have $\|\phi(\mu)\|_{\mathcal{H}}^{2}=\langle\int k(x,\cdot)\mathrm{d}\mu(x),\int k(y,\cdot)\mathrm{d}\mu(y)\rangle_{\mathcal{H}}=\iint k(x,y)\mathrm{d}\mu(x)\mathrm{d}\mu(y)\leq M_{k}$. By Lemma A.3, we then have
@@
with $C_{N,k}:=18NM_{k}$. As in the proof of Corollary 3.8, under our assumptions, we have that $\mathrm{MMD}_{k}\leq CW_{2}$. we obtain that
$\frac{1}{N}\|G_{\mu}^{\mathrm{KME}}-G_{\nu_{K}}^{\mathrm{KME}}\|_{F}^{2}$ $\leq\frac{C_{N,k}}{N}\sum\limits_{i=1}^{N}\|\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\|^{2}_{\mathcal{H}}=\frac{C_{N,k}}{N}\sum\limits_{i=1}^{N}\mathrm{MMD}^{2}_{k}(\mu^{(i)},\nu^{(i)}_{K})$
$\leq\frac{C_{N,k}}{N}\sum\limits_{i=1}^{N}C^{2}W_{2}^{2}(\mu^{(i)},\nu^{(i)}_{K})$
$=C_{N,k}\ \varepsilon_{K},$
if we redefine $C_{N,k}=18NM_{k}C^{2}$.
(i) In order to prove inequality (4.3) for the LOT embedding $\phi:\mu\in\mathcal{P}(\mathcal{X})\mapsto T_{\rho}^{\mu}-\mathrm{id}\in L^{2}(\rho)$ with a.c. reference measure $\rho$, we first note that for any measure $\mu\in\mathcal{P}(\mathcal{X})$, $\|\phi(\mu)\|^{2}_{L^{2}(\rho)}=\int|T_{\rho}^{\mu}(x)-x|^{2}\mathrm{d}\rho(x)\leq\int\mathrm{diam}(\mathcal{X})^{2}\mathrm{d}\rho(x)=\mathrm{diam}(\mathcal{X})^{2}$. Then we can apply Lemma A.3 with constant $C:=C_{N,\mathcal{X}}=18N\mathrm{diam}(\mathcal{X})^{2}$ in (A.12) and finally obtain
$\frac{1}{N}\|G_{\mu}^{\mathrm{LOT}}-G_{\nu_{K}}^{\mathrm{LOT}}\|_{F}^{2}$ $\leq\frac{C_{N,\mathcal{X}}}{N}\sum\limits_{i=1}^{N}\|T_{\rho}^{\mu^{(i)}}-T_{\rho}^{\nu^{(i)}_{K}}\|_{L^{2}(\rho)}^{2}.$ (A.16)
@@
$\|T_{\rho}^{\mu}-T_{\rho}^{\nu}\|_{L^{2}(\rho)}\leq 2\sqrt{\mathrm{diam}(\mathcal{X})L}W_{1}(\mu,\nu)^{1/2}.$ (A.17)
Finally, assembling the inequalities (A.16) and (A.17), and because $W_{1}\leq W_{2}$, we obtain
@@
$=4C_{N,\mathcal{X}}\mathrm{diam}(\mathcal{X})L\frac{1}{N}\sqrt{\left(\sum\limits_{i=1}^{N}W_{2}(\mu^{(i)},\nu^{(i)}_{K})\right)^{2}}$
@@
$=4C_{N,\mathcal{X}}\mathrm{diam}(\mathcal{X})L\frac{1}{N}\sqrt{N\cdot N\varepsilon_{K}}$
$=4C_{N,\mathcal{X}}\mathrm{diam}(\mathcal{X})L\cdot\sqrt{\varepsilon_{K}}$
$=C_{N,\mathcal{X},L}\sqrt{\varepsilon_{K}},$
which concludes the proof. ∎
Proof of Proposition 4.7.
We start by following the classical guidelines in M-estimation in statistics by noticing that
$\mathcal{E}^{PCA}_{q}$ $=E(P_{K}^{\leq q})-E(P^{\leq q})$
$\leq E(P_{K}^{\leq q})+E_{K}(P^{\leq q})-E_{K}(P_{K}^{\leq q})-E(P^{\leq q})$ (A.18)
$\leq 2\sup\limits_{P\in\mathcal{P}_{q}}\bigl{|}E(P)-E_{K}(P)\bigr{|}$
where (A.18) is due to $E_{K}(P^{\leq q})-E_{K}(P_{K}^{\leq q})\geq 0$ as $P_{K}^{\leq q}$ is a minimizer for $E_{K}$.
Now, let us note $\mathbb{T}=\frac{1}{N}\sum\limits_{i=1}^{N}\delta_{\overline{T}^{(i)}}$ and $\mathbb{T}_{K}=\frac{1}{N}\sum\limits_{i=1}^{N}\delta_{\overline{T}^{(i)}_{K}}$. We have that:
$\Bigl{|}E(P)-E_{K}(P)\Bigr{|}$ $=\Bigl{|}-\langle\Sigma,P\rangle_{HS}-(-\langle\Sigma_{K},P\rangle_{HS})\Bigr{|}$
$=\Bigl{|}\frac{1}{N}\sum\limits_{i=1}^{N}\bigl{\langle}\overline{T}^{(i)}\otimes\overline{T}^{(i)},P\bigr{\rangle}_{HS}-\frac{1}{N}\sum\limits_{i=1}^{N}\langle\overline{T}^{(i)}_{K}\otimes\overline{T}^{(i)}_{K},P\bigr{\rangle}_{HS}\Bigr{|}$
$=\Bigl{|}\int\langle T\otimes T,P\rangle_{HS}\mathrm{d}\mathbb{T}(T)-\int\langle T\otimes T,P\rangle_{HS}\mathrm{d}\mathbb{T}_{K}(T)\Bigr{|}$
$=\Bigl{|}\int f_{P}(T)\mathrm{d}(\mathbb{T}-\mathbb{T}_{K})(T)\Bigr{|},$
where $f_{P}:T\in L^{2}(\rho)\rightarrow\langle T\otimes T,P\rangle_{HS}\in\mathbb{R}$. In what follows, we aim to prove that this function $f_{P}$ is Lipschitz, that is
$|f_{P}(T_{1})-f_{P}(T_{2})|$ $=\Bigl{|}\langle T_{1}\otimes T_{1}-T_{2}\otimes T_{2},P\rangle_{HS}\Bigr{|}$
$=\Bigl{|}\langle T_{1}\otimes(T_{1}-T_{2})+(T_{1}-T_{2})\otimes T_{2},P\rangle_{HS}\Bigr{|}$
$\leq\|T_{1}\otimes(T_{1}-T_{2})+(T_{1}-T_{2})\otimes T_{2}\|_{HS}\cdot\|P\|_{HS}$
$\leq\bigl{(}\|T_{1}\|_{L^{2}(\rho)}\cdot\|T_{1}-T_{2}\|_{L^{2}(\rho)}+\|T_{2}\|_{L^{2}(\rho)}\cdot\|T_{1}-T_{2}\|_{L^{2}(\rho)}\bigr{)}\cdot\sqrt{q}$
$=\|T_{1}-T_{2}\|_{L^{2}(\rho)}\bigl{(}\|T_{1}\|_{L^{2}(\rho)}+\|T_{2}\|_{L^{2}(\rho)}\bigr{)}\cdot\sqrt{q}$
$\leq\|T_{1}-T_{2}\|_{L^{2}(\rho)}\cdot 2R\sqrt{q}$
where $R=\max\limits_{x\in\mathcal{X}}\|x\|$. The functional $\frac{f_{P}}{2R\sqrt{q}}$ is therefore 1-Lipschitz. We have that:
$\Bigl{|}E(P)-E_{K}(P)\Bigr{|}$ $=\Bigl{|}\int f_{P}(T)\mathrm{d}(\mathbb{T}-\mathbb{T}_{K})(T)\Bigr{|}$
$=2R\sqrt{q}\Bigl{|}\int\frac{f_{p}(T)}{2R\sqrt{q}}\mathrm{d}(\mathbb{T}-\mathbb{T}_{K})(T)\Bigr{|}$
$\leq 2R\sqrt{q}\Bigl{|}\sup\limits_{f\mbox{ is }1-\text{Lipschitz}}\int f(T)\mathrm{d}(\mathbb{T}-\mathbb{T}_{K})(T)\Bigr{|}$
$\leq 2R\sqrt{q}\cdot\mathcal{W}_{1}(\mathbb{T},\mathbb{T}_{K})$ (A.19)
$=2R\sqrt{q}\min\limits_{P~{}\text{coupling}}\sum\limits_{i,j=1}^{N}P_{ij}\|\overline{T}^{(i)}-\overline{T}^{(j)}_{K}\|_{L^{2}(\rho)}$
$\leq 2R\sqrt{q}\frac{1}{N}\sum\limits_{i=1}^{N}\|T^{(i)}-T^{(i)}_{K}\|_{L^{2}(\rho)}$
$\leq 2R\sqrt{q}\cdot\frac{1}{N}\sum\limits_{i=1}^{N}2\sqrt{\mathrm{diam}(\mathcal{X})L}W_{1}^{1/2}(\mu^{(i)},\nu_{K}^{(i)})$ (A.20)
$=4R\sqrt{q\mathrm{diam}(\mathcal{X})L}\cdot\frac{1}{N}\Biggl{(}\Bigl{(}\sum\limits_{i=1}^{N}W_{2}^{1/2}(\mu^{(i)},\nu_{K}^{(i)})\Bigr{)}^{4}\Biggr{)}^{1/4}$
$\leq 4R\sqrt{q\mathrm{diam}(\mathcal{X})L}\cdot\frac{1}{N}\Bigg{(}N^{3}\sum\limits_{i=1}^{N}W_{2}^{2}(\mu^{(i)},\nu_{K}^{(i)})\Biggr{)}^{1/4}$ (A.21)
$=4R\sqrt{q\mathrm{diam}(\mathcal{X})L}\cdot\varepsilon_{K}^{1/4}$
where (A.19) comes from the Kantorovich-Rubinstein formulation of $\mathcal{W}_{1}$ (see e.g. [51]). Inequality (A.20) follows from [17] and (A.21) follows from the inequality $\bigl{(}\sum\limits_{i=1}^{N}a_{i}\bigr{)}^{4}\leq N^{3}\sum\limits_{i=1}^{N}a_{i}^{4}$. We finally obtain that
$\mathcal{E}_{q}^{PCA}$ $\leq 2\sup\limits_{P\in\mathcal{P}_{q}}|E(P)-E_{K}(P)|$
$\leq 8R\sqrt{q\mathrm{diam}(\mathcal{X})L}\varepsilon_{K}^{1/4},$